# BMG-Q: Localized Bipartite Match Graph Attention Q-Learning for Ride-Pooling Order Dispatch

### v1

@@ Abstract
Index Terms:
Ride-Pooling, Order Dispatch, Multi-agent Reinforcement Learning, Graph Neural Networks.
@@ I Introduction
The widespread adoption of mobile communication and Global Positioning System technology has allowed Transportation Network Companies (TNCs) such as Uber, Lyft, and Didi to provide on-demand mobility services on a global scale [1, 2]. Ever since [3], the advantages of flexible and collaborative ride-sharing operations have become increasingly recognized within the transportation research community. In line with this trend, there has been an expanding body of research on operational policies for ride-sharing, including multi-hop ride-sharing [4], the coordination of ride-hailing with public transportation [5], ride-sharing with passenger transfers [6], integration of ride-sharing with parcel delivery [7], and the coordination of autonomous vehicles with conventional vehicles [8]. These advancements are propelled by breakthroughs in deep learning and Multi-Agent Reinforcement Learning (MARL) frameworks [9, 10, 11].
@@
Nevertheless, several hurdles must be overcome to unlock the full potential of MARL for developing effective, scalable, and robust real-time operational strategies for ride-pooling order dispatch. A principal challenge in this context is the complex interdependence in decision-making among vehicles, which leads to an exponential increase in both state and action spaces within large fleets [12]. One approach to address this is by traditional independent learning approaches, such as Independent Q-Learning (IQL) and Independent Proximal Policy Optimization (IPPO) [13, 14], which ignore the interdependence. In the context of ride-pooling, it is common in the existing literature to combine single-agent independent Reinforcement Learning (RL) with bipartite matching. For instance, [15] and [4] proposed to adopt a Deep Q-Network (DQN) for relocating ride-pooling agents and bipartite matching for order dispatch, and later on, extended it to multi-hop ride-sharing and parcel delivery [7]. To improve the transferability and scalability of the framework, [16] introduced additional techniques such as limited-memory upper confidence bound and reward smoothing. Moreover, [5] coordinated ride-hailing with public transit by encoding the decisions of subway stations into the states of tabular temporal difference learning. Similarly, [6] facilitated ride-pooling with passenger transfer. Yet, the practice of merging independent reinforcement learning with bipartite matching, while improving scalability, often overlooks the agents’ complex interdependence during the RL exploitation and training phases. This can lead to significant overestimation of rewards, a critical concern in highly competitive environments such as ride-pooling, where the pronounced interdependence among agents intensifies the issue.
@@
To accommodate the intricate interdependence among agents, several MARL frameworks have been introduced, including state-of-the-art Centralized Training with Decentralized Execution (CTDE) algorithms such as Multi-Agent Deep Deterministic Policy Gradient (MADDPG) [17], Q-mix [18], Q-tran [19], and Multi-Agent Proximal Policy Optimization (MAPPO) [20]. However, these methods are typically applied to much smaller-scale problems. In contrast, large-scale ride-pooling order dispatch involves thousands of agents, rendering these approaches infeasible. To enhance algorithmic scalability while capturing agent interactions, researchers have explored novel concepts such as Mean-Field MARL [21], where agents interact with an average representation of all other agents. However, in the ride-pooling context [22], this average state may not accurately represent agent interdependence as each agent has unique passengers with different itineraries, and summing this up may lead to misleading information for the decision maker. One approach to overcome this limitation is by considering the Attention-based MARL [23, 24], which instead of relying on average state information, allows distinct neighboring agent to has distinct weights that can be flexibly adjusted over time. Nevertheless, the application of Attention-based MARL is limited in small-scale scenarios and single-passenger ride-hailing [23, 24, 25], where the possibility of multiple riders sharing the same ride is not explicitly considered.
@@
  • We propose a novel BMG-Q framework to address multi-agent interactions in MARL within the context of large-scale ride-pooling order dispatch. The propsoed framework leverages the novel localized bipartite match interdependent Markov Decision Proces (MDP) formulation with the Graph Attention Double Deep Q Network (GATDDQN) as backbones. It captures the interdependence among agents and thus leads to more optimal assignment decisions compared to existing works.
@@ II Related Works
Ride-Pooling Order Dispatch. The operational dynamics of ride-pooling have garnered considerable attention due to their promising yet unpredictable real-time demand, as evidenced by various studies[33, 3, 34]. The nature of this uncertainty, coupled with the full potential of ride-pooling systems, introduces complexity into the process of coordinating vehicles with multiple passengers. Effective coordination requires not only addressing the needs of current passengers but also anticipating the needs of future riders, which includes managing new ride requests and those already being served. Initial investigations in this field have considered short-sighted, or myopic, policies that make vehicular assignments based on presently available information [33, 3]. Specifically, [3] notably advances this by introducing the shareability graph, which identifies possible sharing opportunities between new requests and vehicles on standby. They put forward a batch-matching strategy and crafted a sequential method that divides the decision-making process into vehicle routing and passenger assignment tasks. For more efficient real-time operations, [34] reduces the complexity of the matching problem by limiting the process to pairing a single passenger with a vehicle at each time step. More recent advancements in the field have shifted towards a better incorporation of the uncertainties related to future demand into the decision-making processes via methods such as model predictive control[35, 36, 37], approximate dynamic programming[38, 39, 40], and stochastic integer programming[41]. Note that these works are model-based, requiring explicit characterization of system dynamics and/or future uncertainties.
@@
MARL Framework for Ride-Pooling Dispatch. Given the super-human capabilities of RL and MARL showcased in a range of notable achievements [9, 10, 11], the prospect of crafting practical MARL systems for the real-time optimization of ride-sharing dispatch grows increasingly compelling. While some researchers have endeavored to deploy multi-agent reinforcement learning approaches such as Mean-Field MARL [22], Q-mix [42], and Attention-based MARL [25, 43] in ride-sourcing scenarios, these methods continue to grapple with challenges like stability and scalability when it comes to training in large-scale and complex settings. To address the scalability issue in large-scale ride-sharing systems, it is common in the ride-sharing research community to combine single agent RL (or equivalently, Independent RL) with bipartite match. Specifically, [15, 4] propose to adopt DQN for ride-pooling agents’ relocation and bipartite match for order dispatch, and later on extend it into multi-hop ride sharing and parcel delivery [7]. To improve the transferability and scalability of full deployment of the framework in ride hailing, [16] proposes additional techniques such as limited-memory upper confidence bound and reward smoothing. Moreover, [5] coordinates ride-sourcing with public transit through encoding the decision of subway stations into states of tabular temporal difference learning. [8] coordinates autonomous vehicles with conventional vehicles through two-sided deep reinforcement learning. [6] enable ride-pooling with passenger transfer. However, the aforementioned works have yet to adequately address the intricate interdependencies among vehicles in the ride-pooling context while also ensuring scalability for MARL.
@@
Graph-based MARL. As the computational efficiency and representational power of GNN models such as Graph Convolutional Network (GCN) [44], GraphSAGE [45], Graph Attention Network (GAT) [46], and Relational Graph Convolutional Network (RGCN) [47] gain increasing recognition in complex and adaptive representation learning, researchers have begun to investigate the integration of these potent GNN models with MARL. This nascent area of research seeks to tackle a variety of challenges within MARL, such as the complex task of encoding environmental dynamics from the perspective of individual agents, as well as the decomposition of value functions and the nuanced distribution of credit across the collective team [30]. Specifically for coordination games, on the one hand, [26, 27] propose to adopt graph convolution RL and two-stage attention mechanism to learn abstract interplay representation between agents within graph topology. Following this trend, [48] comprehensively utilizes GATs and RGCN to capture both explicit and implicit relations simultaneously among agents. On the other hand, [28, 29] and [49] propose the idea of coordination graph and utilize GATs to factorize the join team value function or team policy to enable coordination behavior among agents. Despite these advancements, these advances have not yet been applied to ride-pooling, a highly complex and large-scale system, where achieving scalability, stability, and robustness concurrently remains a significant challenge.
@@ III-A MDP Formulation for Ride-Pooling Order Dispatch
where the first term is the starting revenue of a vehicle picking up a new passenger; the second term is the revenue based on the distance between between new order’s origin and destinations (denoted as $Dis$); the third term is the cost of the new passenger waiting to be picked up (with waiting time denoted by $Pickup$); the intuition of the fourth and fifth terms is to give a small penalty if the total additional travel time due to ride pooling compared with a direct non-sharing ride-hailing trip (denoted as $add$) is below the threshold time (denoted as $thre$) but give a heavy penalty if the additional time is above the threshold. For the platform as a whole, the total reward $R_{t}$ at time $t$ is the summation of rewards of all $N$ agents at time $t$:
@@ III-B Independent and Homogeneous Assumptions in MDP
However, in ride-pooling, thousands of agents might need to be simultaneously dispatched, which will lead to a prohibitively high dimensional state and action space for the above centralized MDP. Encountering this challenge, it is common in the previous literature to assume agents are independent [15, 4, 50, 16, 5, 6], i.e., each agent’s transition function and reward function has no interdependence with other agents’ actions (note that our proposed method does not require this assumption), which largely reduces the MDP dimensionality and decentralizes the original transition function from $P(S_{t+1}|S_{t},A_{t})$ to $p(s_{t+1}|s_{t},a_{t})$. The independent assumption modifies the centralized MDP in Equation (6) into:
@@ III-C Merging Independent Learning with Bipartite Matching
Since simply adopting independent assumptions will make the system ignore the complexity compounded by the likelihood of agents making concurrent decisions and fall into conflicts of agents accepting the same order, previous research has often employed a hybrid approach that melds independent reinforcement learning’s policy function or value function with a bipartite matching process [15, 4, 50, 16, 5, 6]. Here we adopt the representative integration of ILP with Independent value function like $Q(s,a)$ to illustrate this idea.
@@
In our study, we adopt the common assumption consistent with many existing literature: each vehicle is assigned only one request per time period. This aligns with many established methodologies, as seen in [15, 51, 52, 43]. Note that this assumption does not impose a significant loss of optimality compared to assigning bundled orders to the same vehicle simultaneously [3]. Specifically, in our context, dispatch decisions are made very frequently (e.g., every minute) in a dynamic manner. If the system intends to assign multiple requests to the same vehicle, it can first assign one order at the current time step, and even before the first order is picked up, it can assign another order to the same vehicle in a subsequent time period. This approach can actually be more optimal than assigning two orders to the same vehicle simultaneously. This is because deferring bundling decisions to future time points, when new information may become available, allows the platform greater flexibility to dynamically adjust decisions under uncertainties.
@@ III-D ILPDDQN Benchmark Framework
1:  Simulator Initialization: Episode Order Requirements, Open Street Routing Mmaching (OSRM) Router Model[32], Matching Distance $R_{match}$, Number of Vehicles $N$.
@@
4:     Initialize: Episode Order Requirements, and Number of Vehicles $N$.
@@
10:           Sample $N$ experience tuples $(s,a,r,s^{\prime})$ in $M$ as mini-batch $D$ and use Equation (13) to update $\theta$.
@@ IV Localized Bipartite Match Interdependent MDP and GATDDQN
While the previously introduced framework that combines independent RL with bipartite matching (i.e., ILPDDQN) significantly enhances scalability, it overlooks the intricate interdependence among agents in the MARL exploitation and training process. This oversight can lead to substantial overestimation of rewards, potentially leading to suboptimal solutions, particularly in the highly competitive environment of ride-pooling. Therefore, in this section, we present our novel Graph-based MARL algorithm, termed as GATDDQN, which effectively captures the agent interdependence with localized bipartite matching graph within a large-scale ride-pooling system. This serves as the novel MARL backbone for our BMG-Q framework.
The rest of this section will proceed as follows. We will initiate our discussion by showing how to build upon the previous MDP framework to incorporate localized bipartite matching. We will then review the fundamentals of classical Graph Attention Neural Network techniques. Following this, we will delve into the structure and formulation of our GATDDQN, which is designed to capture agent interdependence through a localized bipartite matching graph.
@@ IV-B Classical Graph Attention Neural Network
However, unlike the previous case of ride-sourcing [21], the challenge intensifies in ride-pooling dispatch, where fully understanding and aggregating the interdependence within the bipartite match graph becomes more complex. In a ride-pooling environment, each agent may have unique passengers with different itineraries. Traditional GCNs that employ average or max aggregation methods [45] to learn interdependencies within a bipartite match graph can yield misleading or inaccurate information for decision-makers. For illustrative purposes, consider the scenario depicted in Figure 2: an empty ride-pooling agent is surrounded by eight others (in dark blue), each evaluating its decision concerning a new order request heading southeast. Each neighboring agent carries passengers destined for various directions: two heading east, two west, two south, and two north. When this empty agent attempts to assess the interdependencies of its neighbors to make informed decisions, simplistic aggregation methods like averaging or maximum can introduce significant inaccuracies. Averaging the directions might falsely suggest that these agents lack specific destinations, effectively dismissing all directional data (Figure 2a). Conversely, using maximum aggregation could distort the representation, focusing only on a single direction and ignoring the diversity of passenger destinations (Figure 2b).
Figure 2: Illustrative examples for different graph aggregation strategies. An empty ride-pooling agent in black is surrounded by eight vehicles in dark blue, each carrying passengers destined for distinct directions as indicated by the blue arrow. A new request arrives and intends to head southeast, as indicated by the black arrow. While GCN aggregators [45] in (a) or (b) can introduce significant inaccuracies, attention-based aggregation like GAT in (c) allows the empty agent to assign highest weights to its east and south neighbors, which are more compatible with the destination of the new order request.
Therefore, after constructing the localized bipartite match graph for each agent, we can employ GATs to enable an unassigned agent to dynamically weigh its neighbors based on the current scenario. For example, it might assign higher relevance to agents heading south and east, aligning more closely with a southeast-bound order request. To this end, we will first review the basic notations and ideas of classical GATs. In particular, GATs [46] are designed to handle data structured as graphs $G=\{V,E\}$, where $V$ represents the nodes (which are agents in our context), and $E$ represents the edges. A single layer of GATs operates by computing a set of transformations and attention coefficients for each node in the graph.
Firstly, for the message layer, each node or agent $i$ in the graph is transformed using a shared linear transformation (e.g., GraphSAGE [45]), parameterized by a weight matrix $W\in\mathbb{R}^{F^{\prime}\times F}$:
@@
where $a:\mathbb{R}^{F^{\prime}}\times\mathbb{R}^{F^{\prime}}\rightarrow\mathbb{R}$ is a learnable shared attention mechanism, $[s_{i}^{\prime}||s_{j}^{\prime}]$ denotes the concatenation of agent $i$ and $j$’s transformed states, and $\sigma$ is a non-linear activation function. Moreover, to stabilize the learning process and enrich model capacity, GATs employ multi-head attention as an aggregation function:

### v2

@@ I Introduction
The widespread adoption of mobile communication and Global Positioning System technology has allowed Transportation Network Companies (TNCs) such as Uber, Lyft, and Didi to provide on-demand mobility services on a global scale [18, 39]. Ever since [3], the advantages of flexible and collaborative ride-sharing operations have become increasingly recognized within the transportation research community. In line with this trend, there has been an expanding body of research on operational policies for ride-sharing, including the coordination of ride-hailing with public transportation [9], ride-sharing with passenger transfers [43], and the coordination of autonomous vehicles with conventional vehicles [47]. These advancements are propelled by breakthroughs in deep learning and Multi-Agent Reinforcement Learning (MARL) frameworks [4, 28].
@@
Nevertheless, several hurdles must be overcome to unlock the full potential of MARL for developing effective, scalable, and robust real-time operational strategies for ride-pooling order dispatch. A principal challenge in this context is the complex interdependence in decision-making among vehicles, which leads to an exponential increase in both state and action spaces within large fleets [36]. One approach to address this is by traditional independent learning approaches, such as Independent Q-Learning (IQL) and Independent Proximal Policy Optimization (IPPO) [7, 37], which ignore the interdependence. In the context of ride-pooling, it is common in the existing literature to combine single-agent independent Reinforcement Learning (RL) with bipartite matching. For instance, [1] proposed to adopt a Deep Q-Network (DQN) for relocating ride-pooling agents and bipartite matching for order dispatch. To improve the transferability and scalability of the framework, [31] introduced additional techniques such as limited-memory upper confidence bound and reward smoothing. Moreover, [9] coordinated ride-hailing with public transit by encoding the decisions of subway stations into the states of tabular temporal difference learning. Similarly, [43] facilitated ride-pooling with passenger transfer. Yet, the practice of merging independent reinforcement learning with bipartite matching, while improving scalability, often overlooks the agents’ complex interdependence during the RL exploitation and training phases. This can lead to significant overestimation of rewards, a critical concern in highly competitive environments such as ride-pooling, where the pronounced interdependence among agents intensifies the issue.
@@
To accommodate the intricate interdependence among agents, several MARL frameworks have been introduced, including state-of-the-art Centralized Training with Decentralized Execution (CTDE) algorithms such as Multi-Agent Deep Deterministic Policy Gradient (MADDPG) [23], Q-mix [30], Q-tran [35], and Multi-Agent Proximal Policy Optimization (MAPPO) [51]. However, these methods are typically applied to much smaller-scale problems. In contrast, large-scale ride-pooling order dispatch involves thousands of agents, rendering these approaches infeasible. To enhance algorithmic scalability while capturing agent interactions, researchers have explored novel concepts such as Mean-Field MARL [49], where agents interact with an average representation of all other agents. However, in the ride-pooling context, this average state may not accurately represent agent interdependence as each agent has unique passengers with different itineraries, and summing this up may lead to misleading information for the decision maker. One approach to overcome this limitation is by considering the Attention-based MARL [15, 17], which instead of relying on average state information, allows distinct neighboring agent to has distinct weights that can be flexibly adjusted over time. Nevertheless, the application of Attention-based MARL is limited in small-scale scenarios and single-passenger ride-hailing [17, 20], where the possibility of multiple riders sharing the same ride is not explicitly considered.
@@
  • We propose a novel BMG-Q framework to address multi-agent interactions in MARL within the context of large-scale ride-pooling order dispatch. The propsoed framework leverages the novel localized bipartite match interdependent Markov Decision Proces (MDP) formulation with the Graph Attention Double Deep Q Network (GATDDQN) as backbones. It captures the interdependence among agents and thus leads to more effective assignment decisions compared to existing works.
@@ II Related Works
Ride-Pooling Order Dispatch. The operational dynamics of ride-pooling have garnered considerable attention due to their promising yet unpredictable real-time demand, as evidenced by various studies[3]. The nature of this uncertainty, coupled with the full potential of ride-pooling systems, introduces complexity into the process of coordinating vehicles with multiple passengers. Effective coordination requires not only addressing the needs of current passengers but also anticipating the needs of future riders, which includes managing new ride requests and those already being served. Initial investigations in this field have considered short-sighted, or myopic, policies that make vehicular assignments based on presently available information [3, 34]. Specifically, [3] notably advances this by introducing the shareability graph, which identifies possible sharing opportunities between new requests and vehicles on standby. They put forward a batch-matching strategy and crafted a sequential method that divides the decision-making process into vehicle routing and passenger assignment tasks. For more efficient real-time operations, [34] reduces the complexity of the matching problem by limiting the process to pairing a single passenger with a vehicle at each time step. More recent advancements in the field have shifted towards a better incorporation of the uncertainties related to future demand into the decision-making processes via methods such as model predictive control[2], approximate dynamic programming[50], and stochastic integer programming[24]. Note that these works are model-based, requiring explicit characterization of system dynamics and/or future uncertainties.
@@
MARL Framework for Ride-Pooling Dispatch. Given the super-human capabilities of RL and MARL showcased in a range of notable achievements [25, 4, 28], the prospect of crafting practical MARL systems for the real-time optimization of ride-sharing dispatch grows increasingly compelling. While some researchers have endeavored to deploy multi-agent reinforcement learning approaches such as Mean-Field MARL [53], Q-mix [6], and Attention-based MARL [20, 8] in ride-sourcing scenarios, these methods continue to grapple with challenges like stability and scalability when it comes to training in large-scale and complex settings. To address the scalability issue in large-scale ride-sharing systems, it is common in the ride-sharing research community to combine single agent RL (or equivalently, Independent RL) with bipartite match. Specifically, [1] propose to adopt DQN for ride-pooling agents’ relocation and bipartite match for order dispatch. To improve the transferability and scalability of full deployment of the framework in ride hailing, [31] proposes additional techniques such as limited-memory upper confidence bound and reward smoothing. Moreover, [9] coordinates ride-sourcing with public transit through encoding the decision of subway stations into states of tabular temporal difference learning. [47] coordinates autonomous vehicles with conventional vehicles through two-sided deep reinforcement learning. [43] enable ride-pooling with passenger transfer. However, the aforementioned works have yet to adequately address the intricate interdependencies among vehicles in the ride-pooling context while also ensuring scalability for MARL.
@@
Graph-based MARL. As the computational efficiency and representational power of GNN models such as Graph Convolutional Network (GCN) [19], GraphSAGE [13], and Graph Attention Network (GAT) [42] gain increasing recognition in complex and adaptive representation learning, researchers have begun to investigate the integration of these potent GNN models with MARL. This nascent area of research seeks to tackle a variety of challenges within MARL, such as the complex task of encoding environmental dynamics from the perspective of individual agents, as well as the decomposition of value functions and the nuanced distribution of credit across the collective team [26]. Specifically for coordination games, on the one hand, [16, 22] propose to adopt graph convolution RL and two-stage attention mechanism to learn abstract interplay representation between agents within graph topology. [45] propose the idea of coordination graph and utilize GATs to factorize the join team value function or team policy to enable coordination behavior among agents. Despite these advancements, these advances have not yet been applied to ride-pooling, a highly complex and large-scale system, where achieving scalability, stability, and robustness concurrently remains a significant challenge.
@@ III-A MDP Formulation for Ride-Pooling Order Dispatch
where the first term is the starting revenue of a vehicle picking up a new passenger; the second term is the revenue based on the distance between between new order’s origin and destinations (denoted as $Dis$); the third term is the cost of the new passenger waiting to be picked up (with waiting time denoted by $Pickup$); the intuition of the fourth and fifth terms is to give a small penalty if the total additional travel time due to ride pooling compared with a direct non-sharing ride-hailing trip (denoted as addadd) is below the threshold time (denoted as threthre) but give a heavy penalty if the additional time is above the threshold. For the platform as a whole, the total reward $R_{t}$ at time $t$ is the summation of rewards of all $N$ agents at time $t$:
@@ III-B Independent and Homogeneous Assumptions in MDP
However, in ride-pooling, thousands of agents might need to be simultaneously dispatched, which will lead to a prohibitively high dimensional state and action space for the above centralized MDP. Encountering this challenge, it is common in the previous literature to assume agents are independent [1, 31, 9, 43], i.e., each agent’s transition function and reward function has no interdependence with other agents’ actions (note that our proposed method does not require this assumption), which largely reduces the MDP dimensionality and decentralizes the original transition function from $P(S_{t+1}|S_{t},A_{t})$ to $p(s_{t+1}|s_{t},a_{t})$. The independent assumption modifies the centralized MDP in Equation (6) into:
@@ III-C Merging Independent Learning with Bipartite Matching
Since simply adopting independent assumptions will make the system ignore the complexity compounded by the likelihood of agents making concurrent decisions and fall into conflicts of agents accepting the same order, previous research has often employed a hybrid approach that melds independent reinforcement learning’s policy function or value function with a bipartite matching process [1, 31, 9, 43]. Here we adopt the representative integration of ILP with Independent value function like $Q(s,a)$ to illustrate this idea.
@@
In our study, we adopt the common assumption consistent with many existing literature: each vehicle is assigned only one request per time period. This aligns with many established methodologies, as seen in [1, 12, 21, 8]. Note that this assumption does not impose a significant loss of optimality compared to assigning bundled orders to the same vehicle simultaneously [3]. Specifically, in our context, dispatch decisions are made very frequently (e.g., every minute) in a dynamic manner. If the system intends to assign multiple requests to the same vehicle, it can first assign one order at the current time step, and even before the first order is picked up, it can assign another order to the same vehicle in a subsequent time period. This approach can actually be more efficient and effective than assigning two orders to the same vehicle simultaneously. This is because deferring bundling decisions to future time points, when new information may become available, allows the platform greater flexibility to dynamically adjust decisions under uncertainties.
@@ III-D ILPDDQN Benchmark Framework
1: Simulator Initialization: Episode Order Requirements, Open Street Routing Mmaching (OSRM) Router Model[27], Matching Distance $R_{match}$, Number of Vehicles NN.
@@
4:  Initialize: Episode Order Requirements, and Number of Vehicles NN.
@@
10:    Sample NN experience tuples $(s,a,r,s^{\prime})$ in $M$ as mini-batch $D$ and use Equation (13) to update $\theta$.
@@ IV Localized Bipartite Match Interdependent MDP and GATDDQN
While the previously introduced framework that combines independent RL with bipartite matching (i.e., ILPDDQN) significantly enhances scalability, it overlooks the intricate interdependence among agents in the MARL exploitation and training process. This oversight can lead to substantial overestimation of rewards, potentially leading to suboptimal solutions, particularly in the highly competitive environment of ride-pooling. Therefore, in this section, we present our novel Graph-based MARL algorithm, termed as GATDDQN, which effectively captures the agent interdependence with localized bipartite matching graph within a large-scale ride-pooling system. The rest of this section will proceed as follows. We will initiate our discussion by showing how to build upon the previous MDP framework to incorporate localized bipartite matching. We will then review the fundamentals of classical Graph Attention Neural Network techniques. Following this, we will delve into the structure and formulation of our GATDDQN, which is designed to capture agent interdependence through a localized bipartite matching graph.
@@ IV-B Classical Graph Attention Neural Network
However, unlike the previous case of ride-sourcing [49], the challenge intensifies in ride-pooling dispatch, where fully understanding and aggregating the interdependence within the bipartite match graph becomes more complex. In a ride-pooling environment, each agent may have unique passengers with different itineraries. Traditional GCNs that employ average or max aggregation methods [13] to learn interdependencies within a bipartite match graph can yield misleading or inaccurate information for decision-makers. Therefore, after constructing the localized bipartite match graph for each agent, we propose to employ GATs to enable an unassigned agent to dynamically weigh its neighbors based on the current scenario.
To this end, we will first review the basic notations and ideas of classical GATs. In particular, GATs [42] are designed to handle data structured as graphs $G=\{V,E\}$, where $V$ represents the nodes (which are agents in our context), and $E$ represents the edges. A single layer of GATs operates by computing a set of transformations and attention coefficients for each node in the graph. Firstly, for the message layer, each node or agent $i$ in the graph is transformed using a shared linear transformation (e.g., GraphSAGE [13]), parameterized by a weight matrix $W\in\mathbb{R}^{F^{\prime}\times F}$:
@@
where $a:\mathbb{R}^{F^{\prime}}\times\mathbb{R}^{F^{\prime}}\rightarrow\mathbb{R}$ is a learnable shared attention mechanism, $s_{i}^{\prime}||s_{j}^{\prime}]$ denotes the concatenation of agent $i$ and $j$’s transformed states, and $\sigma$ is a non-linear activation function. Moreover, to stabilize the learning process and enrich model capacity, GATs employ multi-head attention as an aggregation function: